# Proportionality for Constrained Public Decisions

### v1

@@ 4 Justified Representation with Cohesiveness
A natural candidate for a pricing type is the standard pricing of unconstrained MES where the price for every issue-decision pair $(a_{t},d)$ is set to $\lambda(a_{t},d)=n$. And with this pricing, that we refer to as unit pricing $\lambda_{\text{unit}}$, we can show that $\text{MES}_{\mathcal{C}}$ satisfies $\text{cohEJR}_{\mathcal{C}}$-1 for unconstrained, binary elections.
@@
For binary election instances, when $\mathcal{C}=\{0,1\}^{m}$, $\text{MES}_{\mathcal{C}}$ with unit pricing $\lambda_{\text{unit}}$ satisfies $\text{cohEJR}_{\mathcal{C}}$-1.
@@
Take an outcome $\boldsymbol{w}$ returned by $\text{MES}_{\mathcal{C}}$ with unit pricing $\lambda_{\text{unit}}$ and consider a $T$-cohesive group of voters $N^{\prime}$. Let us assume that for every voter $i\in N^{\prime}$, it holds that $u_{i}(\boldsymbol{w})<|T|-1$ and then set $\ell=|T|-1$. So to conclude the run of $\text{MES}_{\mathcal{C}}$, each voter in $N^{\prime}$ paid for at most $\ell-1=|T|-2$ issues.
@@
$m-(\ell-1)\frac{m}{\ell+1}=\frac{2m}{\ell+1}=\frac{2m}{|T|}\geqslant\frac{2n}{|N^{\prime}|}.$
The last step follows from the group $N^{\prime}$ being . So now we know that the voters in $N^{\prime}$ hold at least $2n$ in funds when some at the end of $\text{MES}_{\mathcal{C}}$’s run. Thus, we know that at least two issues have not been funded and for at least one of these two issues, at least half of $N^{\prime}$ agree on the decision of this issue (as the election instance is binary) and they hold enough funds to pay for it, hence, we have a contradiction to $\text{MES}_{\mathcal{C}}$ terminating.
Now, assume that some voter $i$ in $N^{\prime}$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$ for a decision on an issue. Since we know that at the end of $\text{MES}_{\mathcal{C}}$’s execution, each voter in $N^{\prime}$ paid for at most $\ell-1=|T|-2$ issues, then at the round $r$ that voter $i$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$ for an issue’s decision, the voters in $N^{\prime}$ collectively held at least $2n$ in funds. Since at least two issues in were not funded, there exists some issue that could have been paid for in round $r$, where voters each pay $\nicefrac{{m}}{{(\ell+1)}}$, contradicting the fact that voter $i$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$ in round $r$. So, we have that this group of voters $N^{\prime}$ cannot exist and that $\text{MES}_{\mathcal{C}}$ satisfies $\text{cohEJR}_{\mathcal{C}}$-1. ∎
This result provides us with an axiom ‘close to’ EJR that we know is always satisfiable when the issues have size-two domains and there are no constraints.
@@ 5 Justified Representation with Agreement
Given the mostly negative results regarding the cohesiveness-EJR notion, we move on to justified representation based on agreement. We justify this move as the notion based on agreement is weaker and yields more positive results in the unconstrained setting. Thus, by assessing it here, we are able to establish a baseline of what can be achieved in terms of EJR-like proportionality guarantees in our constrained model. First, we formalise agreement-based EJR with the following axiom.
@@
Given an election $(\boldsymbol{B},\mathcal{C})$, an outcome $\boldsymbol{w}$ provides $\text{agrEJR}_{\mathcal{C}}$ if for every $T$-agreeing group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w})$-deviation for some $S\subseteq T$ with $|S|\leqslant|T|\cdot\nicefrac{{|N^{\prime}|}}{{n}}$, there exists a voter $i\in N^{\prime}$ such that $u_{i}(\boldsymbol{w})\geqslant\nicefrac{{|N^{\prime}|}}{{n}}\cdot|T|$.
Now, in more unfortunate news, we find that $\text{agrEJR}_{\mathcal{C}}$ is not always satisfiable in general. In fact, the counterexample of Proposition 6 suffices to show this as each voter requires at least $1$ in satisfaction for to $\text{agrEJR}_{\mathcal{C}}$ to be satisfied.
Proposition 13.
There exists an election instance where no outcome provides $\text{agrEJR}_{\mathcal{C}}$ (even when the NFD property holds for $\mathcal{C}$).
We now focus on a particular class of constraints as we import agreement-EJR into our setting. Specifically, we consider a class that allows us to talk about how restrictive, and thus how costly, the fixing of a particular issue-decision pair is.
Akin to work by Rey et al. (2023, 2020), we consider constraints $\mathcal{C}$ that can be equivalently expressed as a set of implications $\textit{Imp}_{\mathcal{C}}$, where each implication in $\textit{Imp}_{\mathcal{C}}$ is a propositional formula with the following form: $\ell_{(a_{x},d_{x})}\rightarrow\ell_{(a_{y},d_{y})}$. This class of constraints allows us, for instance, to express simple dependencies and conflicts such as ‘selecting $x$ means that we must select $y$’ and ‘selecting $x$ means that $y$ cannot be selected’, respectively. These constraints correspond to propositional logic formulas in 2CNF.
@@
Take a set of issues =$\mathcal{I}=\{a,b,c,d,e\}$ for a binary election instance. Here is an example of an implication set:
  • $\textit{Imp}_{\mathcal{C}}=\{(a,1)\rightarrow(b,1),(c,1)\rightarrow(e,0),(d,1)\rightarrow(e,0)\}$. Here, accepting $a$ means that $b$ must also be accepted while accepting either $c$ or $d$ requires the rejection of $e$.
@@
    Add the edge $((a_{y},d_{y}^{*}),(a_{x},d_{x}^{*}))$ for all $d_{y}^{*}\neq d_{y}\in D_{y},d_{x}^{*}\neq d_{x}\in D_{x}$ to $E$ if $\ell_{(a_{x},d_{x})}\rightarrow\ell_{(a_{y},d_{y})}\in\textit{Imp}_{\mathcal{C}}$ and $w_{x}=d_{x}$.
Given such a graph $G_{\boldsymbol{w}}$ for an outcome $\boldsymbol{w}$, we use $G_{\boldsymbol{w}}(a_{x},d_{x})$ to denote the set of all vertices that belong to some path in $G_{\boldsymbol{w}}$ having vertex $(a_{x},d_{x})$ as the source (note that $G_{\boldsymbol{w}}(a_{x},d_{x})$ excludes $(a_{x},d_{x})$).
@@
Consider a binary election instance and take a set of issues =$\mathcal{I}=\{a_{1},a_{2},a_{3},a_{4}\}$ and the implication set $\textit{Imp}_{\mathcal{C}}=\{(a_{1},1)\rightarrow(a_{2},1),(a_{1},1)\rightarrow(a_{3},1),(a_{2},1)\rightarrow(a_{4},1)\}$ of some constraint $\mathcal{C}$. Consider the outcome implication graph for $\boldsymbol{w}_{1}=(0,0,0,0)$ (vertices with no adjacent edges are omitted for readability):
@@
Then, we have $G_{\boldsymbol{w}}(a_{1},1)=\{(a_{2},1),(a_{3},1),(a_{4},1)\}$ and therefore $|G_{\boldsymbol{w}}(a_{1},1)|=3$.
@@
Intuitively, with a $k$-restrictive constraint, if one were to fix/change an outcome $\boldsymbol{w}$’s decision for one issue, this would require fixing/changing $\boldsymbol{w}$’s decisions on at most $k-1$ other issues. So intuitively, when dealing with $k$-restrictive constraints, we can quantify (at least loosely speaking) how ‘difficult’ it is to satisfy a constraint via the use of this value $k$. Thus, we can use this value $k$ to account for the constraint’s difficulty when designing proportionality axioms.
Before assessing how $k$-restrictive constraints affect our goal of providing proportionality, we touch on the computational complexity of checking, for some constraint $\mathcal{C}$, whether there exists a set of implications $\textit{Imp}_{\mathcal{C}}$ that is equivalent to $\mathcal{C}$. For the case of binary elections, this problem been studied under the name of Inverse Satisfiablility and it has been shown that for formulas in 2CNF, the problem is in $\P$ (Kavvadias and Sideri, 1998). So in the remainder of the paper, when we refer to a $k$-restrictive constraint $\mathcal{C}$, we thus assume that $\mathcal{C}$ is expressible using an implication set $\textit{Imp}_{\mathcal{C}}$.
We now import the agreement-EJR notion and an approximate variant into our framework with constraints.
@@
Given an election $(\boldsymbol{B},\mathcal{C})$, some $\alpha\in(0,1]$ and some positive integer $\beta$, an outcome $\boldsymbol{w}$ provides $\alpha\text{-}{\text{agrEJR}_{\mathcal{C}}}\text{-}\beta$ if for every $T$-agreeing group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w})$-deviation for some $S\subseteq T$ with $|S|\leqslant|T|\cdot\nicefrac{{|N^{\prime}|}}{{n}}$, there exists a voter $i\in N^{\prime}$ such that $u_{i}(\boldsymbol{w})\geqslant\alpha\cdot\nicefrac{{|N^{\prime}|}}{{n}}\cdot|T|-\beta$.
With this axiom, we formalise agreement-EJR to our constrained public-decision model with the presence of the multiplicative and additive factors allowing us to measure how well rules satisfy this notion even if they fall short providing the ideal representation. [^3] Observe that we include the axiom’s size requirement on the set $S$ such that a group has an $(S,\boldsymbol{w})$-deviation in order to prohibit considering cases such as a single voter only having an $(S,\boldsymbol{w})$-deviation for $S=\mathcal{I}$ while not intuitively being entitled to that much representation. Note that for the sake of readability, when we have either $\alpha=1$ or $\beta=0$, we omit them from the notation when referring to $\alpha\text{-}{\text{agrEJR}_{\mathcal{C}}}\text{-}\beta$.
@@
Suppose there are four issues =$\mathcal{I}=\{a_{1},a_{2},a_{3},a_{4}\}$ and consider a constraint $\mathcal{C}=\{(1,1,0,0),(1,1,1,0)\}$. Then suppose there are two voters with ballots $\boldsymbol{b}_{1}=(1,1,1,1)$ and $\boldsymbol{b}_{2}=(0,0,0,0)$ so each voter deserves at least $2$ in satisfaction according to agreement-EJR. See that outcome $\boldsymbol{w}=(1,1,0,0)$ provides $\text{agrEJR}_{\mathcal{C}}$ while the outcome $\boldsymbol{w}^{\prime}=(1,1,1,0)$ only provides $\nicefrac{{1}}{{2}}\text{-}{\text{agrEJR}_{\mathcal{C}}}$ as voter $2$ only obtains $1$ in satisfaction whilst having a sufficiently small $(S,\boldsymbol{w})$-deviation for the issue $a_{3}$ (deviating to outcome $\boldsymbol{w}$).
@@
Now, assume that some voter $i\in N^{\prime}$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$ for fixing an issue’s decision. Since we know that at the end of $\text{MES}_{\mathcal{C}}$’s run, each voter in $N^{\prime}$ paid for at most $\ell-1$ issues, then at the round $r$ that voter $i$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$, the voters group $N^{\prime}$ collectively held at least $2nk$ in funds. Since at least $k$ issues in were not funded, there exists some issue that could have been paid for in round $r$, where voters each pay $\nicefrac{{m}}{{(\ell+1)}}$. This contradicts the fact that voter $i$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$ in round $r$. So, we have that this group of voters $N^{\prime}$ cannot exist which concludes the proof. ∎
Towards an even more positive result, and one where we are not limited to binary election instances, we now provide an adaptation of the MeCorA rule (Skowron and Górecki, 2022) . In the unconstrained public-decision model, MeCorA is presented by Skowron and Górecki (2022) as an auction-style variant of MES that allows voter groups to change the decision of an issue all while increasing the price for any further change to this issue’s decision. In our constrained model, groups are allowed to pay for changes to the decisions on sets of issues, as long as these changes represent a feasible deviation.
@@
  • $(iii)$ has an $(S,\boldsymbol{w})$-deviation.
The rule then works as follows. Given a current outcome $\boldsymbol{w}$, it computes, for every non-empty $S\subseteq\mathcal{I}$, the smallest possible value $\rho_{(t,S)}$ for each issue $a_{t}\in S$ such that for some $N^{\prime}$, if voters in $N^{\prime}$ each pay $\rho_{S}=\sum_{a_{t}\in S}\rho_{(t,S)}$ (or their remaining budget), then $N^{\prime}$ is able to ‘update’ the decisions on every $a_{t}\in S$ as per conditions $(i)-(iii)$. If there exists no such voter group for issues $S$ then it sets $\rho_{S}=\infty$.
@@
    updates the current outcome $\boldsymbol{w}$’s decisions on issues in $S$ to the decisions agreed upon by the voters with the associated $(S,\boldsymbol{w})$-deviation,

### v2

@@ 4 Justified Representation with Cohesiveness
A natural candidate for a pricing type is the standard pricing of unconstrained MES where the price for every issue-decision pair $(a_{t},d)$ is set to $\lambda(a_{t},d)=n$. We refer this pricing as unit pricing $\lambda_{\text{unit}}$.
Example 6.
Suppose there are three issues $\mathcal{I}=\{a_{1},a_{2},a_{3}\}$ with domains $D_{t}=\{d_{1},d_{2},d_{3}\}$ for all $t\in\{1,2,3\}$. Then take the constraint to be $\mathcal{C}=\{(d_{1},d_{2},d_{2}),(d_{1},d_{3},d_{3}),(d_{3},d_{2},d_{3})\}$. Then, suppose there are three voters $N=\{1,2,3\}$ with ballots $\boldsymbol{b}_{1}=(d_{1},d_{2},d_{1})$, $\boldsymbol{b}_{2}=(d_{1},d_{2},d_{2})$ and $\boldsymbol{b}_{3}=(d_{3},d_{3},d_{2})$. We consider the execution of $\text{MES}_{\mathcal{C}}$ with a unit pricing $\lambda_{\text{unit}}$ so we have that each issue costs $n=3$ and each voter has a budget of $m=3$. In every round, setting any decision would cost voter $3$ has a price of $3$ as no other voter agrees with them on any decision. Now, take the first round. Voter group $\{1,2\}$ would set either issue $a_{1}$ to decision $d_{1}$, or issue $a_{2}$ to decision $d_{2}$, at a cost of $\nicefrac{{3}}{{2}}$ to each of them. Suppose it was the latter and thus, outcome $(d_{1},d_{3},d_{3})$ is no longer feasible. In the second round, voter $3$ setting either issue $a_{1}$ or issue $a_{3}$ to decision $d_{3}$ would be feasible with respect to the constraint, however, the rule favours voter group $\{1,2\}$ setting issue $a_{1}$ to $d_{1}$ as they (again) pay $\nicefrac{{3}}{{2}}$ each. Thus, outcome $(d_{3},d_{2},d_{3})$ is also not feasible. As voter $3$ does not approve of a decision that remains feasible and the voters in $\{1,2\}$ have spent the of their budgets, the rule terminates with issue $a_{3}$ not being set and the partial outcome $(d_{1},d_{2},\_)$ is returned. However, note that with $(d_{1},d_{2},d_{2})$ there is a feasible completion of this partial outcome in $\mathcal{C}$.
Next, we show that in unconstrained elections, we can use $\text{MES}_{\mathcal{C}}$ to provides fairness guarantees that are ‘close to’ $\text{cohEJR}_{\mathcal{C}}$-1. We first define this approximate version of $\text{cohEJR}_{\mathcal{C}}$.
Definition 8 ($\alpha\text{-}{\text{cohEJR}_{\mathcal{C}}}\text{-}\beta$).
Given an election $(\boldsymbol{B},\mathcal{C})$, some $\alpha\in(0,1]$ and some positive integer $\beta$, an outcome $\boldsymbol{w}$ provides $\alpha\text{-}{\text{cohEJR}_{\mathcal{C}}}\text{-}\beta$ if for every $T$-cohesive group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation for some non-empty $S\subseteq T$, there exists a voter $i\in N^{\prime}$ such that:
$u_{i}(\boldsymbol{w})\geqslant\alpha\cdot|T|-\beta.$
The multiplicative and additive factors in this axiom allow us to measure how well rules satisfy this notion even if they fall short of providing the ideal representation. Now, here is our result for $\text{MES}_{\mathcal{C}}$.
@@
For election instances, when $\mathcal{C}=\{0,1\}^{m}$, $\text{MES}_{\mathcal{C}}$ with unit pricing $\lambda_{\text{unit}}$ satisfies $\nicefrac{{2}}{{(\max_{t\in[m]}|D_{t}|)}}\text{-}{\text{cohEJR}_{\mathcal{C}}}\text{-}1$.
@@
Take an outcome $\boldsymbol{w}$ returned by $\text{MES}_{\mathcal{C}}$ with unit pricing $\lambda_{\text{unit}}$ and consider a $T$-cohesive group of voters $N^{\prime}$. Let us assume that for every voter $i\in N^{\prime}$, it holds that $u_{i}(\boldsymbol{w})<(\nicefrac{{2}}{{\max_{t\in[m]}|D_{t}|}})\cdot|T|-1$ and then set $\ell=(\nicefrac{{2}}{{\max_{t\in[m]}|D_{t}|}})\cdot|T|-1$. So to conclude the run of $\text{MES}_{\mathcal{C}}$, each voter in $N^{\prime}$ paid for at most $\ell-1=(\nicefrac{{2}}{{\max_{t\in[m]}|D_{t}|}})\cdot|T|-2$ issues.
@@
$m-(\ell-1)\frac{m}{\ell+1}=\frac{2m}{\ell+1}=\frac{(\max_{t\in[m]}|D_{t}|)m}{|T|}\geqslant\frac{(\max_{t\in[m]}|D_{t}|)n}{|N^{\prime}|}.$
The last step follows from the group $N^{\prime}$ being $T$-cohesive (and thus, $|N^{\prime}|\geqslant|T|\cdot\nicefrac{{n}}{{m}}$). So now we know that the voters in $N^{\prime}$ hold at least $(\max_{t\in[m]}|D_{t}|)n$ in funds at the end of $\text{MES}_{\mathcal{C}}$’s run. Thus, we know that at least $\max_{t\in[m]}|D_{t}|$ issues have not been funded and for at least one of these issues, at least an $\nicefrac{{1}}{{(\max_{t\in[m]}|D_{t}|)}}$-fraction of $N^{\prime}$ agree on the decision of this issue (as the election instance has at most $\max_{t\in[m]}|D_{t}|$ alternatives for any issue) and they hold enough funds to pay its price of $n$ (given by the unit pricing of $\text{MES}_{\mathcal{C}}$), hence, we have a contradiction to $\text{MES}_{\mathcal{C}}$ terminating.
Now, assume that some voter $i$ in $N^{\prime}$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$ for a decision on an issue. Since we know that at the end of $\text{MES}_{\mathcal{C}}$’s execution, each voter in $N^{\prime}$ paid for at most $\ell-1=(\nicefrac{{2}}{{\max_{t\in[m]}|D_{t}|}})\cdot|T|-2$ issues, it must be that in the round $r$ that voter $i$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$ for an issue’s decision, the voters in $N^{\prime}$ collectively held at least $(\max_{t\in[m]}|D_{t}|)n$ in funds. But, there are at least $\max_{t\in[m]}|D_{t}|$ issues in $\mathcal{I}$ that were not funded, so there exists some issue that could have been paid for in round $r$ where voters each pay $\nicefrac{{m}}{{(\ell+1)}}$. This contradicts the fact that voter $i$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$ in round $r$. So, we have that this group of voters $N^{\prime}$ cannot exist and that $\text{MES}_{\mathcal{C}}$ satisfies $\nicefrac{{2}}{{(\max_{t\in[m]}|D_{t}|)}}\text{-}{\text{cohEJR}_{\mathcal{C}}}\text{-}1$. ∎
Note that for unconstrained binary elections, the above result states that $\text{MES}_{\mathcal{C}}$ with unit pricing $\lambda_{\text{unit}}$ provides $\text{cohEJR}_{\mathcal{C}}$-1.444Note that Skowron and Górecki (2022) gave a similar result for $\text{MES}_{\mathcal{C}}$ with unit pricing $\lambda_{\text{unit}}$ by showing that it satisfies an axiom based on the agreement-EJR notion in the unconstrained binary setting (see Theorem $2$ in their paper). However, the proportionality guarantee weakens when more than two issues are considered for issues in the election. Recall that results by Chandak et al. (2024) showed that we cannot get $\text{cohEJR}_{\mathcal{C}}$ even in the unconstrained setting. Thus, the above positive result shows the promise of $\text{MES}_{\mathcal{C}}$ in providing strong representation guarantees when we have constraints.
@@ 5 Justified Representation with Agreement
Given the mostly negative results regarding the (stronger) cohesiveness-EJR notion, we move on to justified representation based on agreement, which leads to weaker requirements in the unconstrained setting. First, we formalise agreement-based EJR with the following axiom.
@@
Given an election $(\boldsymbol{B},\mathcal{C})$, an outcome $\boldsymbol{w}$ provides $\text{agrEJR}_{\mathcal{C}}$ if for every $T$-agreeing group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation for some $S\subseteq T$ with $|S|\leqslant|T|\cdot\nicefrac{{|N^{\prime}|}}{{n}}$, there exists a voter $i\in N^{\prime}$ such that:
$u_{i}(\boldsymbol{w})\geqslant|T|\cdot\frac{|N^{\prime}|}{n}.$
Intuitively, $\text{agrEJR}_{\mathcal{C}}$ states that a (member of a) $T$-agreeing voter group deserves at least $|T|\cdot\nicefrac{{|N^{\prime}|}}{{n}}$ in satisfaction from an outcome $\boldsymbol{w}$. However, if this is not the case, then fairness is only violated if said voter group can find a suitable set $S\subseteq T$ of issues that they can change towards deviating to a different, feasible outcome. Note a size requirement for the set $S$ has been set to $|S|\leqslant|T|\cdot\nicefrac{{|N^{\prime}|}}{{n}}$. This is made to prevent scenarios where a voter group has an $(S,\boldsymbol{w},\mathcal{C})$-deviation that is ‘too large’, e.g., a voter group that agrees on all issues but constitutes $50\%$ of the population, cannot claim a violation of $\text{agrEJR}_{\mathcal{C}}$ via an $(S,\boldsymbol{w},\mathcal{C})$-deviation where $S=\mathcal{I}$.
@@
Suppose there are three issues $\mathcal{I}=\{a_{1},a_{2},a_{3}\}$ with domains $D_{t}=\{d_{1},d_{2},d_{3}\}$ for all $t\in\{1,2,3\}$. Then take the constraint to be $\mathcal{C}=\{(d_{1},d_{2},d_{3}),(d_{2},d_{2},d_{3})\}$. Then, suppose there are three voters $N=\{1,2,3\}$ with ballots $\boldsymbol{b}_{1}=(d_{1},d_{1},d_{1})$, $\boldsymbol{b}_{2}=(d_{1},d_{1},d_{2})$ and $\boldsymbol{b}_{3}=(d_{3},d_{3},d_{3})$. Take the outcome to be $\boldsymbol{w}=(d_{2},d_{2},d_{3})$. Observe that voter group $\{1,2\}$ agree on the first two issues and they represent a $\nicefrac{{2}}{{3}}$-rd fraction of the population. Since this group has an $(S,\boldsymbol{w},\mathcal{C})$-deviation for $S=\{1\}$ that is of the appropriate size (at most $2\cdot\nicefrac{{2}}{{3}}$), we see that $\boldsymbol{w}=(d_{2},d_{2},d_{3})$ violates $\text{agrEJR}_{\mathcal{C}}$. On the other hand, the outcome $\boldsymbol{w}=(d_{1},d_{2},d_{3})$ provides $\text{agrEJR}_{\mathcal{C}}$.
Unfortunately, we find that $\text{agrEJR}_{\mathcal{C}}$ is not always satisfiable in general. This follows from the counterexample of Proposition 6, as each voter requires at least $1$ in satisfaction for to $\text{agrEJR}_{\mathcal{C}}$ to be satisfied.
Corollary 13.
There exists an election instance where no outcome provides $\text{agrEJR}_{\mathcal{C}}$ (even when $\mathcal{C}$ satisfies NFD).
We now introduce a particular class of constraints that allows us to precisely define how restrictive, and thus how costly, the fixing of a particular issue-decision pair is. Similarly to work by Rey et al. (2020, 2023), we consider constraints $\mathcal{C}$ that can be equivalently expressed as a set of implications $\textit{Imp}_{\mathcal{C}}$, where each implication in $\textit{Imp}_{\mathcal{C}}$ is a propositional formula with the following form where $\ell_{(a_{x},d_{x})}$ is a literal associated with the issue-decision pair $(a_{x},d_{x})$: $\ell_{(a_{x},d_{x})}\rightarrow\ell_{(a_{y},d_{y})}$. This class of constraints allows us, for instance, to express simple dependencies and conflicts such as ‘selecting $x$ means that we must select $y$’ and ‘selecting $x$ means that $y$ cannot be selected’, respectively. These constraints correspond to propositional logic formulas in 2CNF.
Example 8.
Take a set of issues $\mathcal{I}=\{a,b,c,d,e\}$ for a binary election instance. An example of an implication set is $\textit{Imp}_{\mathcal{C}}=\{(a,1)\rightarrow(b,1),(c,1)\rightarrow(e,0),(d,1)\rightarrow(e,0)\}$. Here, accepting issue $a$ means that issue $b$ must also be accepted while accepting either issues $c$ or $d$ requires the rejection of issue $e$.
@@
    If there exists an implication $\ell_{(a_{x},d_{x})}\rightarrow\ell_{(a_{y},d_{y})}\in\textit{Imp}_{\mathcal{C}}$ while $w_{x}=d_{x}$ holds, then add the edge $((a_{y},d_{y}^{*}),(a_{x},d_{x}^{*}))$ for all $d_{y}^{*}\neq d_{y}\in D_{y},d_{x}^{*}\neq d_{x}\in D_{x}$ to $E$.
Given such a graph $G_{\boldsymbol{w}}$ for an outcome $\boldsymbol{w}$, we use $G_{\boldsymbol{w}}(a_{x},d_{x})$ to denote the set of all vertices that belong to some path in $G_{\boldsymbol{w}}$ having vertex $(a_{x},d_{x})$ as the source. Note that $G_{\boldsymbol{w}}(a_{x},d_{x})$ excludes $(a_{x},d_{x})$.
@@
Consider a binary election instance and take a set of issues $\mathcal{I}=\{a_{1},a_{2},a_{3},a_{4}\}$ and the implication set $\textit{Imp}_{\mathcal{C}}=\{(a_{1},1)\rightarrow(a_{2},1),(a_{1},1)\rightarrow(a_{3},1),(a_{2},1)\rightarrow(a_{4},1)\}$ of some constraint $\mathcal{C}$. Consider the outcome implication graph for $\boldsymbol{w}_{1}=(0,0,0,0)$ (vertices with no adjacent edges are omitted for readability):
@@
Then, we have $G_{\boldsymbol{w}_{1}}(a_{1},1)=\{(a_{2},1),(a_{3},1),(a_{4},1)\}$ and therefore $|G_{\boldsymbol{w}_{1}}(a_{1},1)|=3$. Now, consider the outcome implication graph for $\boldsymbol{w}_{2}=(0,1,0,0)$:
$(a_{1},1)$$(a_{3},1)$$(a_{2},0)$$(a_{4},0)$
Here, we have that $G_{\boldsymbol{w}_{2}}(a_{4},0)=\{(a_{2},0)\}$ but note that there is no edge from $(a_{2},0)$ to $(a_{1},0)$ as the latter already holds in the outcome.
@@
Intuitively, with a $k$-restrictive constraint, if one were to fix/change an outcome $\boldsymbol{w}$’s decision for one issue, this would require fixing/changing $\boldsymbol{w}$’s decisions on at most k1k-1 other issues. In the remainder of the paper, when we refer to a $k$-restrictive constraint $\mathcal{C}$, we assume that $\mathcal{C}$ is expressible using an implication set $\textit{Imp}_{\mathcal{C}}$. We also mention in passing that the computational complexity of checking, for some constraint $\mathcal{C}$, whether there exists a set of implications $\textit{Imp}_{\mathcal{C}}$ that is equivalent to $\mathcal{C}$, is a problem that corresponds to Inverse Satisfiablility, which has been shown to be polynomial for formulas in 2CNF (Kavvadias and Sideri, 1998).
We now present an approximate variant of the agreement-EJR notion for public decisions with constraints in a similar manner as we did with $\text{cohEJR}_{\mathcal{C}}$.
@@
Given an election $(\boldsymbol{B},\mathcal{C})$, some $\alpha\in(0,1]$ and some positive integer $\beta$, an outcome $\boldsymbol{w}$ provides $\alpha\text{-}{\text{agrEJR}_{\mathcal{C}}}\text{-}\beta$ if for every $T$-agreeing group of voters $N^{\prime}\subseteq N$ for some $T\subseteq\mathcal{I}$ with an $(S,\boldsymbol{w},\mathcal{C})$-deviation for some non-empty $S\subseteq T$ with $|S|\leqslant|T|\cdot\nicefrac{{|N^{\prime}|}}{{n}}$, there exists a voter $i\in N^{\prime}$ such that:
$u_{i}(\boldsymbol{w})\geqslant\alpha\cdot|T|\cdot\frac{|N^{\prime}|}{n}-\beta.$
Note again that we place a size requirement on the set $S$ on which a group has an $(S,\boldsymbol{w},\mathcal{C})$-deviation so that we rule out cases such as a single voter only having an $(S,\boldsymbol{w},\mathcal{C})$-deviation for $S=\mathcal{I}$ while not intuitively being entitled to that much representation. Note that for readability, when we have $\alpha=1$ or $\beta=0$, we omit them from the notation when referring to $\alpha\text{-}{\text{agrEJR}_{\mathcal{C}}}\text{-}\beta$.
@@
Suppose there is a binary election instance with four issues $\mathcal{I}=\{a_{1},a_{2},a_{3},a_{4}\}$ and consider a constraint $\mathcal{C}=\{(1,1,0,0),(1,1,1,0)\}$. Then suppose there are two voters with ballots $\boldsymbol{b}_{1}=(1,1,1,1)$ and $\boldsymbol{b}_{2}=(0,0,0,0)$ so each voter deserves at least $2$ in satisfaction according to the agreement-EJR notion. See that outcome $\boldsymbol{w}=(1,1,0,0)$ provides $\text{agrEJR}_{\mathcal{C}}$ while the outcome $\boldsymbol{w}^{\prime}=(1,1,1,0)$ only provides $\nicefrac{{1}}{{2}}\text{-}{\text{agrEJR}_{\mathcal{C}}}$ as voter $2$ only obtains $1$ in satisfaction whilst having a sufficiently small $(S,\boldsymbol{w},\mathcal{C})$-deviation for the issue $a_{3}$ (deviating to outcome $\boldsymbol{w}$).
@@
Now, assume that some voter $i\in N^{\prime}$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$ for fixing an issue’s decision. Since we know that at the end of $\text{MES}_{\mathcal{C}}$’s run, each voter in $N^{\prime}$ paid for at most $\ell-1$ issues, then at the round $r$ that voter $i$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$, the voters group $N^{\prime}$ collectively held at least $2nk$ in funds. Since at least $k$ issues in were not funded, there exists some issue that could have been paid for in round $r$, where voters each pay $\nicefrac{{m}}{{(\ell+1)}}$. This contradicts the fact that voter $i$ paid more than $\nicefrac{{m}}{{(\ell+1)}}$ in round $r$. Hence group of voters $N^{\prime}$ cannot exist, concluding the proof. ∎
We now provide a definition of a constraint-aware version of the MeCorA rule (Skowron and Górecki, 2022), in an effort to relax the assumption of working in a binary election instance. In the unconstrained public-decision model, MeCorA is presented as an auction-style variant of MES that allows voter groups to change the decision of an issue all while increasing the price for any further change to this issue’s decision. In our constrained version of this rule, voter groups are allowed to pay for changes to the decisions on sets of issues, as long as these changes represent a feasible deviation.
@@
  • $(iii)$ has an $(S,\boldsymbol{w},\mathcal{C})$-deviation.
The rule then works as follows. Given a current outcome $\boldsymbol{w}$, it computes, for every non-empty $S\subseteq\mathcal{I}$, the smallest possible value $\rho_{(t,S)}$ for each issue $a_{t}\in S$ such that for some $N^{\prime}$, if voters in $N^{\prime}$ each pay $\rho_{S}=\sum_{a_{t}\in S}\rho_{(t,S)}$ (or their remaining budget), then $N^{\prime}$ is able to ‘update’ the decisions on every $a_{t}\in S$ as per conditions (i)(iii)(i)-(iii). If there exists no such voter group for issues $S$ then it sets $\rho_{S}=\infty$.
@@
    updates the current outcome $\boldsymbol{w}$’s decisions on issues in $S$ to the decisions agreed upon by the voters with the associated $(S,\boldsymbol{w},\mathcal{C})$-deviation,